# Improving Steering and Verification in AI-Assisted Data Analysis with Interactive Task Decomposition

### v1

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24); Oct 13–16, 2024; Pittsburgh, PA, USAdoi: XXXXXXX.XXXXXXXbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Natural language interfacesccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Recent advancements in AI and particularly the natural language processing and code generation capabilities of Large Language Models (LLMs) have shown promise to facilitate data science tasks. Specifically, chain-of-thought prompting (Wei et al., 2022) and ReAct prompting (Yao et al., 2022) have emerged as implementation techniques for task decomposition, generating “reasoning” traces, followed by “acting” traces where the LLM can invoke external agents. In data science, such agents might read from datasets and execute code. For example, ChatGPT Data Analysis supports uploading of CSV files, following which the language model and its agents can be used to clean data, display visualizations, answer questions about data, and generate code through its iterative chat interface (OpenAI, 2023).
However, a study of data scientists interacting with ChatGPT without code execution functionality, found that participants were not aware of assumptions that the AI was making, found it tedious to verify the correctness of results, were overwhelmed with long responses, and could not effectively steer the AI when it made mistakes (Chopra et al., 2023b). To better understand user needs around AI steering and verification, we conducted a formative study of data scientists using an AI tool for exploratory data analysis (Section 3) and identified the need for affordances that display the AI’s chain-of-thought as assumptions, and allow the user to modify the assumptions (at any time, even retroactively).
In this paper, we develop approaches to support steering and verification of AI-assisted data analysis that makes decomposition a focal point in the interface, not just an implementation detail (Section 4). By steering, we mean the process of a user interacting with an AI system to guide its output from an initial state to a desired state. By verification, we mean the process of a user interacting with an AI system to understand its output, check its correctness, and deciding about further refinement. We present two systems. The Phasewise system decomposes the problem into three editable phases (assumptions, plan steps, and code) with increasing levels of specificity. The Stepwise system decomposes the solution and displays one step or subgoal of the task with its editable assumptions at a time, visually similar to a computational notebook. In both systems, the code is editable, and contains other features such as the ability to highlight part of the code to ask questions from the AI in the side-pane.
@@
We conducted a controlled, within-subjects experiment (n=18) comparing the Stepwise and Phasewise tools on task decomposition and prompting strategies that support steering, verification, and the user’s perceive utility of the tools. We also developed a baseline tool, called Conversational, that is similar to ChatGPT. Users reported significantly greater control with the two new systems, and found intervention, correction, and verification easier, compared to the baseline.
@@
  • A formative study that highlights two significant limitations of “conversational” AI tools: steering the AI and verifying its output (Section 3).
@@ 2.1. AI-assisted Data Analysis
Previous work has considered how data transformation scripts can be synthesized from demonstrations (e.g., Wrangler (Guo et al., 2011; Kandel et al., 2011)), following an influential line of research that synthesizes programs by example in data wrangling contexts (e.g., (Gulwani, 2011)), which may include natural language (Gulwani and Marron, 2014). These can be constrained to use specific APIs such as pandas, using generator-based synthesis (e.g., AutoPandas (Bavishi et al., 2019)). Scripts can also be synthesized based on heuristics of data quality improvement (e.g, CoWrangler (Chopra et al., 2023a)), and data preparation heuristics can also be learned from corpora (e.g., Auto-Suggest (Yan and He, 2020)).
More recently, a number of commercialized LLM supported data analysis tools have become available. These enable data scientists to access AI-powered chat assistants within their notebook (such as Anaconda (Anaconda, 2024), Databricks (Databricks Assistant, 2024), and Jupyter AI (Jupyter AI, 2024)), and other alternate data-science environments (e.g., DataChat AI (DataChat, 2024), SQL and file editors for Databricks, etc.). The semantic abilities of LLMs, coupled with a chat interface, allows conversational interaction with data, follow-up questions, and highly contextualized responses. Consequently, research has investigated in detail the chatbot paradigm for AI assistance in data analysis and visualization (Hoon et al., 2020; Kassel and Rohs, 2018; Zhi and Metoyer, 2020; Setlur and Tory, 2022; Gu et al., 2023; Dibia, 2023).
@@
Most commercial LLM tools use a simple, turn-based, conversational interaction method that limits steering control and do not provide utilities for verification. Our work expands this design space by introducing novel interaction methods based on decomposing the task into structured and editable assumptions, actions, and other components.
@@ 2.2. Verifying LLM Outputs and their Reliability
Gu et al. (Gu et al., 2024) created a design probe similar to ChatGPT Data Analyst (OpenAI, 2023) with an added sidebar for inspecting intermediary variables. They conducted a study with 22 participants using the design probe to understand common behaviors for verifying the AI’s response to a natural language query and dataset. They found two main behaviors within the verification workflow: procedural-oriented and data-oriented which in many cases were tightly coupled and participants frequently switched between an intermediary variable and the code that outputted it. In contrast with how data analysts verify their work in any tool-assisted (non-AI) data analysis workflows, there is now a much bigger demand for verification when users “offload” an entire data analysis task to an LLM.
In our work, we explicitly ask the AI to show its assumptions and reasoning in a structured (and editable) way so that users could focus on them and make decisions. We also include features such as “side queries” that allow users to pose exploratory questions, build up assumptions, and then add those assumptions to the AI generation workflow.
@@ 2.3. Steering LLMs
Research on the metacognitive demands of generative AI identifies decomposition and structured generation as potential aids (Tankelevitch et al., 2024). Suh et al. (Suh et al., 2023) explore hierarchical text generation at different abstraction levels to assist with sensemaking and managing information overload from large text quantities. They also introduce structured generation (Suh et al., 2024), where user’s prompt is first used to generate dimensions that make the model’s responses vary, and then responses are generated according to those dimensions. Additionally, Masson et al. (Masson et al., 2024) propose principles of direct manipulation for LLMs: continuous representation of objects of interest, physical actions to localize prompt effects, and reusable prompts.
@@
However, these approaches hide the AI’s reasoning and decomposition process, leaving users without insight into the “how” and “why” behind the generated code. The decision to manually infer the AI’s reasoning from the output and coming up with explicit actions to edit and refine the grounded utterances is left entirely to the user. While this might be an acceptable trade-off in systems that generate short programs (e.g. typical spreadsheet formulas or SQL queries), it is unclear how this approach would extend to longer and more complex data analysis scripts.
@@ 3. Formative Study
To understand the challenges involved in data analysis with conversational AI assistants, we conducted a formative study with 15 participants (12 male, 3 female). We used the Noteable plugin for ChatGPT, providing a web-based computational notebook for code execution. Users could input natural language (NL queries into ChatGPT, which would generate and send code to Noteable. Upon execution, Noteable returned the results to ChatGPT for further analysis and follow-up queries. At the time of study, ChatGPT with the Noteable plugin was the only available tool offering features similar to those in ChatGPT Data Analysis (formerly Code Interpreter).
@@ 3.1. Results
Participants struggled with the linear conversational format of interaction with ChatGPT, often requesting an “undo button” for AI’s misunderstandings or errors. Without this, they tried workarounds, asking the AI to ‘‘undo the last step’’ (F9) or to ‘‘ignore the previous data cleansing steps and do it from scratch’’ (F5).
@@ 3.2. Design Goals and Rationale
Based on our formative study, relevant prior work, and how such AI tools use Chain-of-Thought prompting for task decomposition and execution, we identified the following design goals to improve user and control over the AI-assisted data analysis process.
DG.1 The system should visually distinguish assumptions that the AI is making, and contrast them from their corresponding actions.
DG.2 The system should allow users to edit, modify, and update the AI’s assumptions and actions.
DG.3 The system should provide various intervention points for the user to understand and edit the AI’s reasoning.
DG.4 The system should facilitate users in efficiently retrieving essential information to validate responses and to inform decision-making at each intervention point.
Currently, LLM-based tools such as ChatGPT Plugins and Advanced Data Analysis employ strategies like Chain-of-Thought prompting (Wei et al., 2022) and ReAct prompting (Yao et al., 2022). These methods decompose the input task into natural language ”reasoning” traces (e.g. dataset and task assumptions), and subsequently ”act” using code execution agents that receive code as an input (which the LLM is capable of generating) and return the output.
Despite these advancements in task decomposition, a disconnect persists between users and the AI’s reasoning, as evidenced by the frequent attempts to steer and correct the AI. This disconnect underscores the necessity for our design goals to focus on making task decomposition more interactive and transparent. Participants had to identify the assumptions behind mistakes and then attempt to correct them using follow-up prompts. This disrupted the data analysis workflow and was not always effective, leading to frustration. Thus, the first three design goals aim to enhance the user’s ability to visually distinguish, directly edit, and intervene in the AI’s reasoning process.
Furthermore, participants often engaged in iterative prompting to explore the dataset or intermediary results to build up assumptions and steer the AI accordingly. Prior work has also showed that a lack of comprehension of the code and the exact data operations impacts data analysts’ confidence in their verification (Gu et al., 2024). However, due to the linearity of the conversational interface, users needed to mix clarification and exploratory prompts with their actual analytical tasks. This resulted in a workflow that was slow, prone to inaccuracies, and disruptive to the primary task. Additionally, the AI itself can aid the user in forming proper assumptions by retrieving relevant information from the dataset at intervention points. This led to our fourth design goal.
@@ 4. System Design
We address design goals DG.1 and DG.2 through interactive task decomposition. This involves: (1) Prompting the LLM to generate its chain-of-thought reasoning as assumptions and corresponding actions about the input task and dataset; and (2) parsing and rendering LLM output as structured, editable UI components, allowing users to refine the AI’s proposed plan. We refer to these as the editable assumptions throughout the system design.
In addressing DG.3 to provide proper intervention points, we encounter trade-offs in balancing the number, amount of information presented, and the degree of control provided at each intervention point. This led us to develop two alternative approaches. The Phasewise system, which gives users greater control over the entire analysis plan from the outset, but with fewer intervention points, and requires the user to understand more information at each step. Conversely, the Stepwise allows more focused control by decomposing the task into step-by-step subgoals, increasing intervention opportunities, reducing the information overload per step, but with less structured control over the entire task.
Similarly, to address DG.4, we balanced the amount of information displayed for verification and decision making at each intervention point. The Phasewise system aids AI-based information retrieval by displaying relevant dataset columns, allowing inspection of column statistics and AI assumptions, as well as distinguishing required and suggested steps in the execution plan. To support user-led exploration in both systems, we allocated a “sidebar” on the ride side of the screen where users can: (a) select portions of the AI-generated code and ask questions about them; (b) ask natural language queries for data exploration; and (c) generate code from natural language description for the user to manually incorporate into the AI-generated code.
In the following sections we outline the core features shared between the Phasewise and Stepwise systems, and explain how they differ. Lastly, we describe the Conversational system, serving as a baseline similar to ChatGPT’s Advanced Data Analysis plugin for our user evaluation.
@@ 4.1. Core System Features
Figure 2. The starting input for all systems, which includes a button to upload datasets a, selectable datasets to be included in the analysis b, and a text input for entering the natural language description that specifies the data analysis task c.
Figure 3. Overview of the Phasewise system’s task flow, which decomposes tasks into three stages. Input + Output Assumptions, allows users to upload a dataset a, manage column-based AI assumptions b, inspect column-based descriptive statistics c, and add columns missed by the AI d. Execution Plan contains the AI’s editable natural language plan for solving the task f, which includes user-selectable optional steps g. Code and Output contains AI generated code for solving the task and includes an code editor h, intermediary variable inspector i, and the code output j. Section 4.1.2 details these features.
@@ 4.1.1. Task Input
Data analysis begins with dataset(s) and a task specification. Dataset Input: Data is loaded using the Input Query interface (Figure 2, a). Users can then select one or more datasets relevant to the task (Figure 2, b). The Python server generates a summary of each selected dataset with sample values for all columns. This summary is passed to the LLM to build its initial set of assumptions for the data analysis task.
Task Specification Input: After selecting relevant datasets, users specify their data analysis task in the text field and press submit to start the data analysis process (Figure 2, c).
@@ 4.1.2. Task Decomposition: Phasewise System
A) Input and Output Assumptions: After loading a dataset (Figure 3, a), this component displays all the columns that the AI found to be relevant to the task, and for each column it displays several editable assumptions regarding the task (Figure 3, b). Assumptions can pertain to data type, uniformity, units, sorting order, etc. Users can delete columns they find unrelated to the task, or add columns that the AI incorrectly did not select (Figure 3, d). Within each column, users can edit, add, or remove assumptions for that column (Figure 3, b). For each column, users can ”inspect” descriptive statistics (Figure 3, c), including a frequency table of sample values for categorical columns. Additionally, the entire dataset can be viewed by clicking on the ”open” button a, with the selected columns highlighted to help the user leverage the columns to build up assumptions. Finally, the task’s output assumptions can be viewed and changed to edit, add, or remove assumptions to steer the final output (Figure 3, e).
B) Execution Plan: Using the assumptions, including edits, the system generates a list of natural language steps for solving the task (Figure 3, f). Steps are editable and the user may add or remove steps. The model is also prompted to include optional steps that are rendered as selectable steps with a checkbox (Figure 3, g). After the user is satisfied with the plan, they can proceed to generating and running the code.
C) Code and Output: Here the AI generates code to solve the task based on the previous two components. The code is immediately executed and displayed in an editor to allow modification and re-execution (Figure 3, h). Users can inspect the dataframe and variables used in the code execution (Figure 3, i), and see the code output (Figure 3, j)

### v2

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: The 37th Annual ACM Symposium on User Interface Software and Technology; October 13–16, 2024; Pittsburgh, PA, USAbooktitle: The 37th Annual ACM Symposium on User Interface Software and Technology (UIST ’24), October 13–16, 2024, Pittsburgh, PA, USAdoi: 10.1145/3654777.3676345isbn: 979-8-4007-0628-8/24/10ccs: Human-centered computing Natural language interfacesccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Recent advancements in AI and particularly the natural language processing and code generation capabilities of Large Language Models (LLMs) have shown promise to facilitate data science tasks. Specifically, chain-of-thought prompting (Wei et al., 2022) and ReAct prompting (Yao et al., 2022) have emerged as implementation techniques for task decomposition, generating “reasoning” traces, followed by “acting” traces where the LLM can invoke external agents. In data science, such agents might read from datasets and execute code. For example, ChatGPT Data Analysis supports uploading CSV files, after which its language model and code execution agents can be used to clean data, display visualizations, and answer questions about data through its iterative chat interface (OpenAI, 2023).
However, a study of data scientists using ChatGPT without code execution functionality found that participants were unaware of the AI’s assumptions when solving the task, found verifying the correctness of results tedious, were overwhelmed by long responses, and could not effectively steer the AI when it made mistakes (Chopra et al., 2023b).
To better understand user behavior, needs, and challenges when performing exploratory data analysis with a conversational AI tool, we conducted a formative study involving 15 participants (Section 3). The study identified steering and verification as the primary limitations of conversational AI tools. Steering refers to the user’s interaction with the AI to guide its output from an initial state to a desired outcome. Verification refers to the user’s interaction to understand the AI’s output, check its correctness, ensure no incorrect assumptions were used, and decide on further refinement. The study also highlighted the need for new affordances that decompose and display the AI’s chain-of-thought reasoning as structured and interactive assumptions, enabling users to modify them at any time, even retroactively.
Based on this core requirement we developed two systems that make “decomposition” a focal point in the interface, not just an implementation detail (Section 4). First, the Phasewise system decomposes the problem into three editable phases (assumptions, planning, and code) with increasing levels of specificity. Second, the Stepwise system decomposes the task into separate steps (visually similar to a computational notebook), displaying editable assumptions and their corresponding code one step at a time until the task is complete.
@@
We conducted a controlled, within-subjects experiment (n=18) comparing the Stepwise and Phasewise tools on task decomposition and prompting strategies that support steering, verification, and the user’s perceived utility of the tools. We also developed a baseline tool, called Conversational, similar to ChatGPT with code execution. Users reported significantly greater control with the two new systems, and found intervention, correction, and verification easier, compared to the baseline.
@@
  • A formative study that identifies the limitations of “conversational” AI tools in terms of steering them and verifying their output.
@@ 2.1. AI-assisted Data Analysis
Previous work has considered how data transformation scripts can be synthesized from demonstrations (e.g., Wrangler (Guo et al., 2011; Kandel et al., 2011)). This follows an influential line of research that synthesizes programs from examples in data wrangling contexts (e.g., (Gulwani, 2011)), which may include natural language (Gulwani and Marron, 2014). These can be constrained to use specific APIs such as pandas, using generator-based synthesis (e.g., AutoPandas (Bavishi et al., 2019)). Scripts can also be synthesized based on heuristics of data quality improvement (e.g, CoWrangler (Chopra et al., 2023a)), and data preparation heuristics can also be learned from corpora (e.g., Auto-Suggest (Yan and He, 2020)).
More recently, a number of commercialized LLM supported data analysis tools have become available. These enable data scientists to access AI-powered chat assistants within their notebook (such as Anaconda (Anaconda, 2024), Databricks (Databricks Assistant, 2024), and Jupyter AI (Jupyter AI, 2024)), and other alternate data-science environments (e.g., DataChat AI (DataChat, 2024), SQL and file editors for Databricks, etc.). The semantic abilities of LLMs, coupled with a chat interface, allows conversational interaction with data, follow-up questions, and highly contextualized responses. Consequently, research has investigated the chatbot paradigm for AI assistance in data analysis and visualization in detail (Hoon et al., 2020; Kassel and Rohs, 2018; Zhi and Metoyer, 2020; Setlur and Tory, 2022; Gu et al., 2023; Dibia, 2023).
@@ 2.2. Verifying LLM Outputs and their Reliability
Furthermore, previous research on AI coding assistants emphasizes that engaging users in verifying the process of how the AI generates code can maximize user experience, efficiency, and the predictability of obtaining a helpful response (Kazemitabaar et al., 2024). For instance, CodeHelp (Liffiton et al., 2023) incorporates a Sufficiency Check step that engages users in refining the AI’s understanding of the task by prompting them to clarify uncertainties or provide missing context.
Gu et al. (Gu et al., 2024) created a design probe similar to ChatGPT Data Analysis (OpenAI, 2023) with an added sidebar for inspecting intermediary variables. They conducted a study with 22 participants using the design probe to understand common behaviors for verifying the AI’s response to a natural language query and dataset. They found two main behaviors within the verification workflow: procedural-oriented and data-oriented which in many cases were tightly coupled and participants frequently switched between an intermediary variable and the code that outputted it. In contrast with how data analysts verify their work in any tool-assisted (non-AI) data analysis workflows, there is now a much bigger demand for verification when users “offload” an entire data analysis task to an LLM.
In our work, we explicitly ask the AI to show its assumptions and reasoning in a structured (and editable) way, paired with their corresponding actions, so that users could focus on them and make decisions. We also include features such as “side queries” that allow users to pose exploratory questions, build up assumptions, and then add those assumptions to the AI generation workflow.
@@ 2.3. Steering LLMs
Currently, most commercial LLM tools (e.g., ChatGPT Data Analysis (OpenAI, 2023) or ChatGPT with Noteable (not, [n. d.])) use a turn-based conversational method, where the AI attempts to solve problems with minimal intervention points. Typically, users can only steer the AI after it has generated an entire solution, using follow-up prompts, which limits steering control. To address such limitations, Masson et al. (Masson et al., 2024) propose principles of direct manipulation (Hutchins et al., 1985) for steering LLMs in other contexts: continuous representation of objects of interest, physical actions to localize prompt effects, and reusable prompts. Furthermore, research on the metacognitive demands of generative AI identifies decomposition and structured generation as potential aids (Tankelevitch et al., 2024). Suh et al. (Suh et al., 2023) explore hierarchical text generation at different abstraction levels to assist with sensemaking and managing information overload from large text quantities. They also introduce structured generation (Suh et al., 2024), where user’s prompt is first used to generate dimensions that make the model’s responses vary, and then responses are generated according to those dimensions.
@@
However, these approaches hide the AI’s reasoning and decomposition process, leaving users without insight into the “how” and “why” behind the generated code. Users are left to manually infer the AI’s reasoning from the output and determine explicit actions to edit and refine the grounded utterances. While this might be an acceptable trade-off in systems that generate short programs (e.g., typical spreadsheet formulas or SQL queries), it is unclear how this approach would extend to longer and more complex data analysis scripts. Our work expands this design space by not only displaying the AI’s assumptions in a structured way but also enable users to directly edit these assumptions as a novel method of steering the AI to control its output.
@@ 3. Formative Study
To explore the challenges of data analysis with conversational AI assistants, we conducted a formative study with 15 participants (12 male, 3 female, 0 non-binary) using the Noteable plugin for ChatGPT (not, [n. d.]). At the time of the study, Noteable was the only publicly available tool offering features similar to ChatGPT Data Analysis (formerly Code Interpreter). With Noteable, participants could upload datasets to a Noteable project and enter a natural language (NL) descriptions of their data analysis task in ChatGPT. In response, ChatGPT would generate code cells in the Noteable project, which Noteable would execute. ChatGPT then displayed Noteable’s output including any tables or visualizations, and generated an interpretation of the results. ChatGPT would then continue generating code if the task was incomplete, or asked users for additional information if required.
@@ 3.1. Results
Furthermore, our results indicate that each participant engaged in linear conversations consisting, on average, 20 AI-generated messages (SD=5). They experienced lengthy responses upon each interaction point, averaging 24 lines of AI-generated code (SD=21, Max=152) and 134 words of the AI’s interpretation of the output (SD=104, Max=717). Participants often lost track of the long conversation history and struggled with finding, verifying, and fixing accumulated assumptions. As a result, they requested an “undo button” to fix accumulated assumptions made by the AI. Without this feature, they tried workarounds, asking the AI to ‘‘undo the last step’’ (F9) or to ‘‘ignore the previous data cleansing steps and do it from scratch’’ (F5).
@@ 3.2. Design Goals and Rationale
Our formative study highlighted steering and verification as the most common user interactions. Based on our findings, along with relevant prior work and how such AI tools use chain-of-thought prompting for task decomposition and execution, we established the following design goals to enhance user control over the AI-assisted data analysis process.
First, participants struggled to understand the AI’s reasoning. They often tried to manually infer the underlying assumptions from its generated code, verify them, and then correct them with follow-up prompts. moThis was evident from the 36 prompts that they used to explicitly fix incorrect assumptions made by the AI. Therefore, DG.1 proposes visually separating each different assumption from its corresponding actions (code) and allowing users to directly edit and update them.
Second, participants were overwhelmed by long responses and lost track of the long conversation history. Consistent with previous studies (Chopra et al., 2023b), DG.2 recommends adding intervention points in the AI’s responses to help users focus on smaller information chunks. Additionally, steering operations should update only relevant sections at each intervention point, rather than adding new outputs to the main thread.
Lastly, participants frequently used prompts for data exploration, result verification, and code explanation. While these were useful, they often derailed the main conversation thread from solving the task. To address this, DG.3 suggests enabling side conversations and other methods to assist users in verifying assumptions without cluttering the main thread.
@@ 4. System Design
We address design goal DG.1 through interactive task decomposition. This involves: (1) prompting the LLM to generate its chain-of-thought reasoning as NL assumptions and corresponding actions about the input task and dataset; and (2) rendering the LLM’s output as structured, editable UI components, allowing users to refine the AI’s proposed plan. We refer to these as editable assumptions that represent the AI’s reasoning based on the task and dataset (e.g., pattern of values in a dataset column.)
In addressing DG.2 to provide proper intervention points, we encounter trade-offs in balancing the number, amount of information presented, and the degree of control provided at each intervention point. This led us to develop two alternative approaches. The Phasewise system, which gives users greater control over the entire analysis plan from the outset, but with fewer intervention points, and requires the user to understand more information at each step. Conversely, the Stepwise allows more focused control by decomposing the task into step-by-step subgoals, increasing intervention opportunities, reducing the information overload per step, but with less structured control over the entire task.
Similarly, to address DG.3, we balanced the amount of information displayed for verification and decision making at each intervention point. The Phasewise system aids AI-based information retrieval by displaying relevant dataset columns, allowing inspection of column statistics and AI assumptions, as well as distinguishing required and suggested steps in the execution plan. To support user-led exploration in both systems, we allocated a “sidebar” on the ride side of the screen where users can: (a) select portions of the AI-generated code and ask questions about them; (b) ask natural language queries for data exploration; and (c) generate code from natural language description for the user to manually incorporate into the AI-generated code.
In the following, sections we outline the core features shared between the Phasewise and Stepwise systems, and explain how they differ. Lastly, we describe the Conversational system, serving as a baseline similar to ChatGPT’s Advanced Data Analysis plugin for our user evaluation.
@@ 4.1. Core System Features
Figure 2. The starting input for all systems, which includes a button to upload datasets Refer to caption, selectable datasets to be included in the analysis Refer to caption, and a text input for entering the natural language description that specifies the data analysis task Refer to caption.
Figure 3. Overview of the Phasewise system’s task flow, which decomposes tasks into three stages. Input + Output Assumptions, allows users to upload a dataset Refer to caption, manage column-based AI assumptions Refer to caption, inspect column-based descriptive statistics Refer to caption, add columns missed by the AI Refer to caption, and edit assumptions about the task’s output Refer to caption. Execution Plan contains the AI’s editable natural language plan for solving the task Refer to caption, which includes user-selectable optional steps Refer to caption. Code and Output contains AI generated code for solving the task and includes an code editor Refer to caption, intermediary variable inspector Refer to caption, and the code output Refer to caption. Section 4.1.2 details these features.
@@ 4.1.1. Task Input
Data analysis begins with dataset(s) and a task specification. Dataset Input: Data is loaded using the Input Query interface (Figure 2, [Uncaptioned image]). Users can then select one or more datasets relevant to the task (Figure 2, [Uncaptioned image]). The Python server generates a summary of each selected dataset with sample values for all columns. This summary is passed to the LLM to build its initial set of assumptions for the data analysis task.
Task Specification Input: After selecting relevant datasets, users specify their data analysis task in the text field and press submit to start the data analysis process (Figure 2, [Uncaptioned image]).
@@ 4.1.2. Task Decomposition: Phasewise System
A) Input and Output Assumptions: After loading a dataset (Figure 3, [Uncaptioned image]), this component displays all the columns that the AI found to be relevant to the task, and for each column it displays several editable assumptions regarding the task (Figure 3, [Uncaptioned image]). Assumptions can pertain to data type, uniformity, units, sorting order, etc. Users can delete columns they find unrelated to the task, or add columns that the AI incorrectly did not select (Figure 3, [Uncaptioned image]). Within each column, users can edit, add, or remove assumptions for that column (Figure 3, [Uncaptioned image]). For each column, users can “inspect” descriptive statistics (Figure 3, [Uncaptioned image]), including a frequency table of sample values for categorical columns. Additionally, the entire dataset can be viewed by clicking on the “open” button [Uncaptioned image], with the selected columns highlighted to help the user leverage the columns to build up assumptions. Finally, the task’s output assumptions can be viewed and changed to edit, add, or remove assumptions to steer the final output (Figure 3, [Uncaptioned image]).
B) Execution Plan: Using the assumptions, including edits, the system generates a list of natural language steps for solving the task (Figure 3, [Uncaptioned image]). Steps are editable and the user may add or remove steps. The model is also prompted to include optional steps that are rendered as selectable steps with a checkbox (Figure 3, [Uncaptioned image]). After the user is satisfied with the plan, they can proceed to generating and running the code.
C) Code and Output: Here the AI generates code to solve the task based on the previous two components. The code is immediately executed and displayed in an editor to allow modification and re-execution (Figure 3, [Uncaptioned image]). Users can inspect the dataframe and variables used in the code execution (Figure 3, [Uncaptioned image]), and see the code output (Figure 3, [Uncaptioned image])